The distance S in metres covered by particle at time T, seconds is given by S= 2t²– 3t² – 5t, find:


(i) it's speed In the 5th second.


(ii) the distance covered in the 3rd second.


(iii) the acceleration at T=5 seconds.

1
Expert's answer
2022-05-24T13:56:09-0400
"s=2t^3-3t^2-5t"

(i)

"v(t)=s'(t)=6t^2-6t-5\\
v(5)=6*5^2-6*5-5=115\:\rm m/s"

(ii)

"d=s(5)-s(4)\\
=(2*5^3-3*5^2-5*5)\\-(2*4^3-3*4^2-4*5)=90\:\rm m"

(iii)

"a=12t-6\\
a(5)=12*5-6=54\:\rm m/s^2"